Write an equation of the line containing the given point and perpendicular to the given line. (8,-6); 3x+5y=8. The equation of the line is y=
Found 2 solutions by richwmiller, nyc_function:
Answer by richwmiller(17219)

Just curious.
Had you looked for similar problems before posting?
Had you looked for lessons before posting?

Answer by nyc_function(2741)

The line we are looking for is perpendicular to the given equation, which means that its slope is the negative reciprocal of the slope of the given equation.
3x + 5y = 8
5y = -3x + 8
y = (-3/5)(x) + 8
The negative reciprocal of (-3/5) is (5/3) because (-3/5) times (5/3) = -1.
We now plug (5/3) and your point (8, -6) into the point-slope formula and solve for y.
y - (-6) = (5/3)(x - 8)
y + 6 = (5/3)(x) - 40/3
y = (5/3)(x) - 40/3 - 6
y = (5/3)(x) - 22/3